CAN SOMEONE PLEASE HELP ME AND SHOW ME HOW TO WORK OUT THESE PROBLEMS.I IF U BREAK IT DOWN FOR ME I WILL UNDERSTAND IT BETTER.
1) A BAG PF DIMES AND QUARTERS IS WORTH $3.55. THERE ARE FOUR MORE THAN DOUBLE THE NUMBER OF DIMES AS QUARTERS. HOW MANY OF THE COINS ARE DIMES?

2)THE PERIMETER OF A TRIANGLE IS 23 INCHES. THE LONGEST SIDE EXCEEDS THE SHORTER BY 7 INCHES WHILE THE THIRD SIDE IS TWICE THE SHORT SIDE. WHAT IS THE LENGTH OF THE LONGEST SIDE?

3) TWICE A NUMBER LESS FIVE IS LESS THAN THE NUMBER LESS THREE. WHAT'S THE VALUE OF THE NUMBER?

4) THE SUM OF TWO NUMBERS IS 25, AND THEIR DIFFERENCE IS 9. WHAT ARE THE TWO NUMBERS?

Answer by stanbon(75887)

1) A BAG OF DIMES AND QUARTERS IS WORTH $3.55. THERE ARE FOUR MORE THAN DOUBLE THE NUMBER OF DIMES AS QUARTERS. HOW MANY OF THE COINS ARE DIMES?
-----
Value Equation: 10d + 25q = 355 cents
Quantity Eq:::: q = 2d+4
----
Substitute for "q" and solve for "d":
10d + 25(2d+4) = 355
10d + 50d + 100 = 355
60d = 255
Comment: d is not a whole number;
The problem is inconsistent.
=====================================
2)THE PERIMETER OF A TRIANGLE IS 23 INCHES. THE LONGEST SIDE EXCEEDS THE SHORTER BY 7 INCHES WHILE THE THIRD SIDE IS TWICE THE SHORT SIDE. WHAT IS THE LENGTH OF THE LONGEST SIDE?
Let the shortest side be "x":
Longest = "x+7"
3rd side = 2x
----
Equation:
x + 2x + x+7 = 23
4x = 16
x = 4 inches (shortest)
2x = 8 inches (3rd side)
x+7 = 11 inches (longest)
=============================
3) TWICE A NUMBER LESS FIVE IS LESS THAN THE NUMBER LESS THREE. WHAT'S THE VALUE OF THE NUMBER?
Equation:
2x-5 < x-3
x < 2
==================

4) THE SUM OF TWO NUMBERS IS 25, AND THEIR DIFFERENCE IS 9. WHAT ARE THE TWO NUMBERS?
Equation:
x + y = 25
x - y = 9
------
Add and solve for "x":
2x = 34
x = 17
Solve for "y":
x+y = 25
17 + y = 25
y = 8
========================
Cheers,
Stan H.
==========================